A small block of mass m is kept at
the left end of a larger block of
mass M and length L
which is on a horizontal surface
such that the system can slide.The
system is started
towards right with an initial velocity
v.The friction coefficient between
the surface and the
larger block is x and the friction
coefficient between the blocks is
x/2.Find the time elapsed
before the smaller block separates
from the bigger block.(sir please
show the steps in details)

Grade:12th pass

1 Answers

Arun Kumar IIT Delhi
9 years ago
Let a1 and a2 be the accelerations of ma and M respectively.
Here, a1 > a2 so that m moves on M
it would be easiest if we see this in relative frame of M
see in relative frame

relative distance to be covered is l wrt to M
so we find a rel to M =a1-a2
and 1/2 a*t^2 = l

and for the force lower body M will have two force friction from the box below =-k(M+m)g
and fromt the upper body -k/2(mg)
a2=-(kmg-2kMg)/2M
a1=-k/2g
take the difference
-kg(M+m)/2M

now apply the kinematics formulae